# Differentially Private Matchings

### v1

@@ 3 Lower Bound for Explicit Matchings
Given a graph $G=(V,E)$, let $\nu(G)$ denote the size of a maximal matching in $G$ (while this is not fully well-defined, all graphs on which we use this definition will have the property that all maximal matchings have the same size, and so are actually maximum matchings). Given a (randomized) algorithm $\mathcal{A}$ which outputs some graph $H=(V,E_{H})$ when run on input graph $G=(V,E)$, let $\nu_{\mathcal{A}}(G)=\nu((V,E\cap E_{H}))$ be a maximal matching in the graph consisting of edges that are in both $G$ and $H$.
We now begin our proof of Theorem 1.1. At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that the marginal probability of choosing each non-edge is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
@@ 4 $\varepsilon$-Local Edge Differentially Private Implicit Matchings
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 1.2, which is restated below for readibility. We modify Algorithm 2 in the later sections for various different models. See 1.2
@@
High Level Overview
@@ 1 $\eta\leftarrow 1/2$
2 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2+(2(1+\eta))/\eta}$
@@ 15                   $u$ releases $A(u)\leftarrow i$
18       if $A(v)>i$ ($v$ has not satisfied its matching condition)  then
@@ 28                   $M(v)\leftarrow M(v)\cup\{u\}$
Algorithm 2 $\varepsilon$-LEDP Maximal Matching
4.2 Privacy Guarantees
@@
By the choice of $\varepsilon^{\prime}=\frac{\varepsilon}{2+(2(1+\eta))/\eta}$ (Algorithm 2), we can conclude that the algorithm is $\varepsilon$-LEDP. ∎
4.3 Utility
We now prove that the implicitly output $b$-matching also contains a maximal matching, with high probability.
@@ 4.3 Utility
For $b=O\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs an implicit $b$-matching that contains a maximal matching in the billboard model, with high probability.
@@
For each iteration $i$, we will show if node $v=v_{i}$ has not satisfied its matching condition and there exists at least $1$ node that comes after it in the order, then it is matched with at least one of the total number of nodes it can be matched with. That is, let $b^{*}(v)\in[0,b]$ denote the maximum number of nodes later than $v$ in the permutation which $v$ can match with (up to $b$); we wish to show that the number of nodes $v$ matches with after iteration $i$ is at least $1$ if $b^{*}(v)\geq 1$. We will then show this suffices to guarantee that our $b$-matching contains a maximal matching. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma 2.18, We have that each of the following holds with probability $1-1/n^{3c}$:
@@
where the last line follows since we know that $b\geq\frac{576c\ln(n)}{\varepsilon}+1$. In other words, $v$ was already matched from a previous iteration.
@@
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$. As a reminder, we are in the case that Algorithm 2 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 2. If the returned $r^{*}>\log_{1+\eta}(k\eta/(8c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\eta}{8c\ln{n}}\right)$. Thus, we have that $p_{r}=\frac{8c\ln(n)}{k\eta}$ so the expected size of $W_{r}(v)$ is $\frac{8c\ln(n)}{\eta}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{8c\ln(n)}{\eta}$. By a multiplicative Chernoff Bound (Theorem 2.19), the true size of $W_{r}(v)$ is at most $16c\log(n)/\eta\leq 16c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
@@
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\eta}{8c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Hence, by our assumption that $k\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$, if $|M(v)|=0$, it holds that $|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Thus, $\mathbb{E}[|W_{r^{*}}(v)|]\geq(1+\eta)^{-1}\cdot(b-36c\ln(n)/\varepsilon^{\prime})\geq 144c\ln(n)/\varepsilon$ and by a Chernoff bound, $|W_{r^{*}}(v)|\geq 1$, with probability at least $1-\frac{1}{n^{4c}}$. It follows that $v$ is matched with at least 1 neighbor at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ and we are done. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
@@
5 $O(\log n)$ Round $\varepsilon$-LEDP Matchings
@@
See 1.4
At a high level, our algorithm performs multiple rounds of matching where in each round some nodes are proposers and others are receivers. A node participates in a round if it has not satisfied its matching condition; a node that participates in the current round is an active node. The proposers are chosen randomly and propose to a set of nodes to match. We call the set of nodes each node proposes to as the proposal set. Receivers are active nodes that are not chosen as proposers and receive the proposers’ proposals. Then, each receiver chooses a subset of proposals to accept. The algorithm continues until all nodes satisfy the matching condition.
@@ 7       Return $r$
5.2 Privacy Guarantees
@@
5.3 Utility and Number of Rounds
We first prove the following lemma which will help prove our utility bounds. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. We first prove the following crucial lemma which gives the “progress” of the algorithm after each round. Namely, in order to ensure we produce a maximal matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their matching capacity every round (hence, would not participate in future matching rounds). For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the below analysis, we call the set of nodes which have not exceeded their matching capacity and have a non-zero number of neighbors which have not reached their matching capacity, the hopeful nodes. A node which is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma which states that a vertex $w$ will choose $r_{w}=0$, with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon$.
@@ 5.3 Utility and Number of Rounds
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$. By Lemma 2.18, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
@@
If $w$ is still active in round $i$, this means that $w$ did not meet its matching condition in round $i-1$. By Lemma 2.18, this means that with probability at least $1-\frac{1}{n^{4c}}$, it holds that $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ because the noise drawn in Algorithm 3 is at least $-12c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$ and the noise drawn in Algorithm 3 is at most $24c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since, $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}<<27c\log(N)/\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}>>\varepsilon^{\prime}$, by our proof for Lemma 5.2, node $w$ will choose a $|W_{r_{w}}(w)|$ can be as large as $50c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since we chose $\eta=0.5$, we will pick an $r_{w}$ where $33c\log(n)/\varepsilon^{\prime}<(50\cdot 2)/3c\log(n)/\varepsilon^{\prime}\leq\mathbb{E}[|W_{r}(w)|]\leq 50c\log(n)/\varepsilon^{\prime}$. Otherwise, if the expectation is less than $33c\log(n)/\varepsilon^{\prime}$, then $w$ will pick all available edges (and hence pick $r_{w}=0$). Now we show the concentration of the size with sufficiently high probability via the Chernoff bound. We can use the Chernoff bound since the edges picked for $W_{r_{w}}(n)$ are i.i.d. chosen at random using coins that are flipped heads with probability $(1+\eta)^{-r_{w}}$. Specifically, by Theorem 2.19, if we take $\psi=0.4$, then, $|W_{r_{w}}(w)|<32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at most $\exp\left(-\frac{0.4^{2}\cdot 33c\log(n)}{3\varepsilon^{\prime}}\right)<\frac{1}{n^{3c}}$. Hence, together, $w$ will pick an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ or it will pick $r_{w}=0$. ∎
@@
We now show that at least a fixed constant fraction of the hopeful, good edges become unhopeful each round with high constant probability. For each good, hopeful edge $e=(u,v)$, oriented from $u$ to $v$, the marginal probability that the source of the edge is a proposer and the sink is a receiver is $1/4$. Let the degree of $v$ in the induced subgraph consisting of all hopeful edges be $\deg(v)$; since $e$ is oriented from $u$ to $v$, it means that $\deg(u)\leq\deg(v)$. The marginal probability that $u$’s proposed set contains $v$ is at least $\min(1,32c\log_{16/15}(n)/(\varepsilon^{\prime}\cdot\deg(u)))\geq\min(1,32c\log_{16/15}(n)/(\varepsilon^{\prime}\cdot\deg(v)))$ by Lemma 5.3. Thus, the probability that $u$ is a proposer, $v$ is a receiver, and $u$ proposes $v$ is at least $\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. Receiver $v$ can either choose to match with $u$ or not. If $v$ does not match with $u$, that means by Lemma 5.3, $v$ picked a set that did not contain $u$ and has reached its matching threshold, with probability at least $1-\frac{1}{n^{3c}}$.
@@
The expected number of edges $v$ receives as part of all proposals is then $\deg(v)\cdot\min(1,32c\log_{16/15}(n)/(4\varepsilon^{\prime}\cdot\deg(u)))\geq\deg(v)\cdot\min(1,32c\log_{16/15}(n)/(4\varepsilon^{\prime}\cdot\deg(v)))\geq\min(\deg(v),32c\log_{16/15}(n)/(4\varepsilon^{\prime}))$. Thus, in expectation, $v$ satisfies its matching condition. Now, we show the concentration on the probability that $v$ satisfies the matching condition. By the Chernoff bound and suppose $\psi=0.4$, the probability that $<10c\log(n)/(4\varepsilon^{\prime})$ proposers propose to $v$ is upper bounded by $\exp\left(\frac{0.4^{2}\cdot 32c\log_{16/15}(n)/\varepsilon^{\prime}}{12}\right)<\exp\left(\frac{2c\log(n)}{12\varepsilon^{\prime}}\right)<\frac{1}{n^{4c}}$. Hence, by the union bound over $n$ nodes, with probability at least $1-\frac{1}{n^{3c}}$, each good receiver satisfies its matching condition. Each hopeful, good edge has $1/4$ probability that it is going from a proposer to a receiver. The expected number of hopeful edges that are good and are going from proposers to receivers is then at least $\frac{H}{8}\cdot(1-1/n^{3})$; these are also the expected number of hopeful edges that become unhopeful. By the Markov bound, the probability that at most $1615H/$ 16 hopeful edges remain hopeful is $\frac{7/8+\frac{1}{n^{3}}}{15/16}=14/15+\frac{15}{16n^{3}}$. This means that with probability at least $1/16$ (assuming $n>10$—if $n\leq 10$, we can run our algorithm from Section 4 in $O(1)$ rounds), at least $H/16$ of the hopeful edges become unhopeful. ∎
@@
Using all of the above lemmas, we prove the final utility guarantee of our algorithm which returns a maximal matching in $O(\log n)$ rounds, with high probability.
Lemma 5.5.
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that contains a maximal matching, with high probability, when $b=\Omega(\log^{2}(n)/\varepsilon)$, in $O(\log(n))$ rounds.
Proof.
@@
By Lemma 5.4, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. Because the Chernoff bound holds for binomials, by Theorem 2.19, if we set $\psi=0.5$, then the probability that less than $16c\log(n)$ rounds are successful is at most $\exp\left(\frac{0.5^{2}\cdot 32c\log(n)}{3}\right)<\frac{1}{n^{8c}}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.

### v2

@@ 3 Lower Bound for Explicit Matchings
Given a graph $G=(V,E)$, let $\nu(G)$ denote the size of a maximal matching in $G$ (while this is not fully well-defined, all graphs on which we use this definition will have the property that all maximal matchings have the same size, and so are actually maximum matchings). Given a (randomized) algorithm $\mathcal{A}$ which outputs some graph $H=(V,E_{H})$ when run on input graph $G=(V,E)$, let $\nu_{\mathcal{A}}(G)=\nu((V,E\cap E_{H}))$ denote the size of a maximal matching in the graph consisting of edges that are in both $G$ and $H$.
At a very high level, we will give a collection of input graphs and argue that any differentially private algorithm must be a poor approximation on at least one of them. This is in some ways similar to the standard “packing arguments” used to prove lower bounds for differentially private algorithms (see, e.g., [DMN23]), but instead of using a packing / averaging argument we will instead use a symmetry argument. We will argue that without loss of generality, any DP algorithm for our class of inputs must be highly symmetric in that the marginal probability of choosing each non-edge is identical. This makes it easy to argue that high-quality algorithms cannot satisfy differential privacy.
3.1 Proof of Theorem 1.1
@@ 4 $\varepsilon$-Local Edge Differentially Private Implicit Matchings
In this section, we state our main algorithm for maximal matching (Algorithm 2) and prove its guarantees in Theorem 1.3, which is restated below for readibility. We modify Algorithm 2 in the later sections for various different models. See 1.3
@@
High-Level Overview
@@ 1 $\eta\leftarrow 1/2$
2 $\varepsilon^{\prime}\leftarrow\frac{\varepsilon}{2(1+2\eta)/\eta}$
@@ 9       $A(u)\leftarrow\infty$
12       for each node $u\in V$ such that $A(u)>i$  do
             // $u$ has not satisfied its matching condition
@@ 15                   $u$ releases $A(u)\leftarrow i$
18       if $A(v)>i$  then
             // $v$ has not satisfied its matching condition
@@ 28                   $M(v)\leftarrow M(v)\cup\{u\}$
Algorithm 2 $\varepsilon$-LEDP Matching
4.2 Proof of Theorem 1.3
The pseudocode for Theorem 1.3 can be found in Algorithm 2. We prove its privacy and approximation guarantees separately. In Section 4.2.1, we first prove that Algorithm 2 is $\varepsilon$-DP. The utility proofs can then be found in Section 4.2.2.
4.2.1 Privacy Guarantees
@@
By the choice of $\varepsilon^{\prime}=\frac{\varepsilon}{2(1+2\eta)/\eta}$ (Algorithm 2), we can conclude that the algorithm is $\varepsilon$-LEDP. ∎
4.2.2 Utility
We first prove that the implicitly output $b$-matching is at least half the size of a maximum matching, with high probability.
@@ 4.2.2 Utility
Let $M\subseteq E$ be an edge subset satisfying the following maximality condition: every vertex is either matched to at least $b^{\prime}$ edges of $M$, or all its unmatched neighbors are matched to at least $b^{\prime}$ edges of $M$. Then if $\operatorname{OPT}$ denotes the size of a maximum $b^{\prime}$-matching, we have
$|M|\geq\frac{1}{2}\operatorname{OPT}\,.$
We will mostly use Lemma 4.2 with $b^{\prime}=1$, but we state and prove the general version needed for Section 10.
@@
Let $M^{*}$ denote an arbitrary $b^{\prime}$-matching. We would like to show that $\frac{|M|}{|M^{*}|}\geq\frac{1}{2}$. Since $\frac{|M\setminus M^{*}|}{|M^{*}\setminus M|}\leq\frac{|M|}{|M^{*}|}$, we may assume without loss of generality that $M\cap M^{*}=\varnothing$ after adjusting the $b^{\prime}$ values per vertex.
Let $S\subseteq V$ denote the set of nodes $v$ with at least $b^{\prime}(v)$ matched edges in $M$. We claim that every edge of $M^{*}$ must have an endpoint in $S$. Otherwise, there is an edge $uv$ such that $u,v$ are matched to less than $b^{\prime}(u),b^{\prime}(v)$ neighbors, respectively, but is not in $M$, contradicting the maximality condition.
Now, each $v\in S$ is incident to at most $b^{\prime}(v)$ edges in $M^{*}$, while every edge of $M$ is incident to at most $2$ edges of $S$. Moreover, each node in $S$ is incident to at least $b^{\prime}(v)$ edges of $M$ by construction. Thus by a double counting argument, $|M^{*}|\leq\sum_{v\in S}b^{\prime}(v)\leq 2|M|$ as desired. ∎
Lemma 4.3.
For $b=O\left(\frac{\log(n)}{\varepsilon}\right)$, Algorithm 2 outputs an implicit $b$-matching that is at least $\nicefrac{{1}}{{2}}$ the size of a maximum matching, with high probability.
Proof.
For each iteration $i$, we will show if node $v=v_{i}$ has not satisfied its matching condition and there exists at least $1$ node that comes after it in the order, then it is matched with at least one of the total number of nodes it can be matched with. That is, let $b^{*}(v)\in[0,b]$ denote the maximum number of nodes later than $v$ in the permutation which $v$ can match with (up to $b$); we wish to show that the number of nodes $v$ matches with after iteration $i$ is at least $1$ if $b^{*}(v)\geq 1$. By Lemma 4.2, this implies the utility guarantee. Finally, we show that with high probability, we do not match with more than $b$ nodes.
First, all of the Laplace noise (Algorithms 2, 2, 2 and 2) are drawn from distributions with expectation 00. Let $X$ be a random variable drawn from one of these Laplace distributions. By Lemma A.1, we have that each of the following holds with probability $1-1/n^{3c}$:
@@
where the first line follows since we know that $b\geq\frac{576c\ln(n)}{\varepsilon}+1$. In other words, $v$ was already matched from a previous iteration.
@@
Case II.b.1: For the remainder of the proof, we assume $k\coloneqq|W_{0}(v)|\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$. As a reminder, we are in the case that Algorithm 2 does execute. Let $r^{*}$ be the $r$ chosen in Algorithm 2. If the returned $r^{*}>\log_{1+\eta}(k\varepsilon^{\prime}/(8c\ln{n}))$, we know that the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\ln(n)/\varepsilon^{\prime}\leq b$ did not hold for $r=\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$. Thus, we have that $p_{r}=(1+\eta)^{r}=\frac{8c\ln(n)}{k\varepsilon^{\prime}}$ so the expected size of $W_{r}(v)$ is $\frac{8c\ln(n)}{\varepsilon^{\prime}}$; that is, $\mathbb{E}[|W_{r}(v)|]=\frac{8c\ln(n)}{\varepsilon^{\prime}}$. By a multiplicative Chernoff Bound (Theorem A.2), the true size of $W_{r}(v)$ is at most $16c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{4c}}$. Since the condition $|\widetilde{M}(v)|+|\widetilde{W}_{r}(v)|+12c\log(n)/\varepsilon^{\prime}\leq b$ didn’t hold for such $r$, we have that
@@
Case II.b.2: If $r^{*}\leq\log_{1+\eta}\left(\frac{k\varepsilon^{\prime}}{8c\ln{n}}\right)$, there are two scenarios. Either $r^{*}=0$, then we match $v$ with all nodes which are still available so clearly the number of nodes $v$ is matched with is at least $b^{*}(v)$, or $r^{*}>1$. In the latter scenario, we know that the threshold is exceeded for $r^{*}-1$, so we have $|\widetilde{M}(v)|+|\widetilde{W}_{r^{*}-1}(v)|+12c\ln(n)/\varepsilon^{\prime}\geq b$. Conditioning on $\mathcal{E}$, $|M(v)|+|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$. Hence, by our assumption that $k\geq\frac{8c\ln(n)}{\varepsilon^{\prime}}$, if $|M(v)|=0$, it holds that $|{W}_{r^{*}-1}(v)|\geq b-36c\ln(n)/\varepsilon^{\prime}$ Thus, by a Chernoff bound, with probability at least $1-\frac{1}{n^{4c}}$
$|W_{r^{*}}(v)|\geq\frac{1-\eta}{1+\eta}\mathbb{E}[|W_{r^{*}-1}(v)|]\geq\frac{1-\eta}{1+\eta}(b-36c\ln(n)/\varepsilon^{\prime})\geq 1\,.$
It follows that $v$ is matched with at least 1 neighbor at the end of this iteration.
Taking a union bound over all $r$ and $v\in V$, all Chernoff bounds hold regardless of the value of $r^{*}$ with probability at least $1-\frac{1}{n^{2c}}$ for $c\geq 3$. Taking into account the conditioning on $\mathcal{E}$, our algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for $c\geq 3$.
@@
5 $O(\log n)$-Round $\varepsilon$-LEDP Implicit Matchings
@@
See 1.6
High-Level Overview
Our algorithm performs multiple rounds of matching where in each round some nodes are proposers and others are receivers. A node participates in a round if it has not satisfied its matching condition; a node that participates in the current round is an active node. The proposers are chosen randomly and propose to a set of nodes to match. We call the set of nodes each node proposes to as the proposal set. Receivers are active nodes that are not chosen as proposers and receive the proposers’ proposals. Then, each receiver chooses a subset of proposals to accept. The algorithm continues until all nodes satisfy the matching condition.
@@ 7       Return $r$
5.2 Proof of Theorem 1.6
The pseudocode for Theorem 1.6 is presented in Algorithm 3. We prove its privacy and utility guarantees separately. We first prove that it satisfies $\varepsilon$-DP in Section 5.2.1 and then prove its approximation guarantees in Section 5.2.2.
5.2.1 Privacy Guarantees
@@
5.2.2 Utility and Number of Rounds
The goal of this section is to prove the following lemma, which gives the formal approximation guarantees and round complexity of Algorithm 3.
Lemma 5.2.
For any $\varepsilon\in(0,1)$, Algorithm 3 returns a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching, with high probability, when $b=\Omega(\log^{2}(n)/\varepsilon)$, in $O(\log(n))$ rounds.
Towards proving Lemma 5.2, we first prove the following lemma. Recall that our algorithm randomly divides the population into (roughly) half proposers and half receivers. Then, the proposers first propose their proposal sets, and then the receivers respond with their match sets for each of $32c\cdot\lceil\log_{16/15}(n)\rceil$ rounds. The following crucial lemma gives the “progress” of the algorithm after each round. Namely, to ensure we produce a maximal matching in $O(\log n)$ rounds, with high probability, roughly a constant fraction of the nodes would need to exceed their matching capacity every round (hence, would not participate in future matching rounds). For convenience in the below proofs, when we write $\log(n)$, we mean $\log_{1+\eta}(n)$.
In the analysis below, we call the set of nodes that have not exceeded their matching capacity and have a non-zero number of neighbors who have not reached their matching capacity, the hopeful nodes. A node that is not hopeful is called unhopeful.
First, note that every active node in any round $i$ is either a proposer or receiver by definition. We first prove the following lemma, which states that a vertex $w$ will choose $r_{w}=0$ with high probability, as a proposer or receiver if its degree is less than $27c\log(n)/\varepsilon$.
@@ 5.2.2 Utility and Number of Rounds
Suppose $X\sim\text{Lap}(4/\varepsilon^{\prime})$ and $Y\sim\text{Lap}(2/\varepsilon^{\prime})$ are the noise random variables chosen in Algorithms 4 and 4, respectively. Since $w$ has induced degree less than $27c\log(n)/\varepsilon^{\prime}$, it will pick $r_{w}=0$ if $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$. By Lemma A.1, $X\leq 24c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$ and $Y\leq 12c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{6c}}$. Then, $X+Y\leq 36c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{5c}}$ when $n\geq 2$. ∎
@@
If $w$ is still active in round $i$, this means that $w$ did not meet its matching condition in round $i-1$. By Lemma A.1, this means that with probability at least $1-\frac{1}{n^{4c}}$, it holds that $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ because the noise drawn in Algorithm 3 is at least $-12c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$ and the noise drawn in Algorithm 3 is at most $24c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since, $|M(w)|\leq 2c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}<<27c\log(N)/\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}>>\varepsilon^{\prime}$, by our proof for Lemma 5.3, node $w$ will choose a $|W_{r_{w}}(w)|$ can be as large as $50c\log(n)/\varepsilon^{\prime}$ with probability at least $1-\frac{1}{n^{3c}}$. Since we chose $\eta=0.5$, we will pick an $r_{w}$ where $33c\log(n)/\varepsilon^{\prime}<(50\cdot 2)/3c\log(n)/\varepsilon^{\prime}\leq\mathbb{E}[|W_{r}(w)|]\leq 50c\log(n)/\varepsilon^{\prime}$. Otherwise, if the expectation is less than $33c\log(n)/\varepsilon^{\prime}$, then $w$ will pick all available edges (and hence pick $r_{w}=0$). Now we show the concentration of the size with sufficiently high probability via the Chernoff bound. We can use the Chernoff bound since the edges picked for $W_{r_{w}}(n)$ are i.i.d. chosen at random using coins that are flipped heads with probability $(1+\eta)^{-r_{w}}$. Specifically, by Theorem A.2, if we take $\psi=0.4$, then, $|W_{r_{w}}(w)|<32c\log_{16/15}(n)/\varepsilon^{\prime}$ with probability at most $\exp\left(-\frac{0.4^{2}\cdot 33c\log(n)}{3\varepsilon^{\prime}}\right)<\frac{1}{n^{3c}}$. Hence, together, $w$ will pick an $r_{w}$ where $|W_{r_{w}}(w)|\geq 32c\log_{16/15}(n)/\varepsilon^{\prime}$ or it will pick $r_{w}=0$. ∎
@@
We now show that at least a fixed constant fraction of the hopeful, good edges become unhopeful each round with high constant probability. For each good, hopeful edge $e=(u,v)$, oriented from $u$ to $v$, the marginal probability that the source of the edge is a proposer and the sink is a receiver is $1/4$. Let the degree of $v$ in the induced subgraph consisting of all hopeful edges be $\deg(v)$; since $e$ is oriented from $u$ to $v$, it means that $\deg(u)\leq\deg(v)$. The marginal probability that $u$’s proposed set contains $v$ is at least
$\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(u)}\right)\geq\min\left(1,\frac{32c\log_{16/15}(n)}{\varepsilon^{\prime}\cdot\deg(v)}\right)$
by Lemma 5.4. Thus, the probability that $u$ is a proposer, $v$ is a receiver, and $u$ proposes $v$ is at least $\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$. Receiver $v$ can either choose to match with $u$ or not. If $v$ does not match with $u$, that means by Lemma 5.4, $v$ picked a set that did not contain $u$ and has reached its matching threshold, with probability at least $1-\frac{1}{n^{3c}}$.
The expected number of edges $v$ receives as part of all proposals is then
$\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(u)}\right)$ $\geq\deg(v)\cdot\min\left(1,\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}\cdot\deg(v)}\right)$
$=\min\left(\deg(v),\frac{32c\log_{16/15}(n)}{4\varepsilon^{\prime}}\right)\,.$
Thus, in expectation, $v$ satisfies its matching condition. Now, we show the concentration on the probability that $v$ satisfies the matching condition. By the Chernoff bound and suppose $\psi=0.4$, the probability that $<10c\log(n)/(4\varepsilon^{\prime})$ proposers propose to $v$ is upper bounded by
$\exp\left(\frac{0.4^{2}\cdot 32c\log_{16/15}(n)/\varepsilon^{\prime}}{12}\right)<\exp\left(\frac{2c\log(n)}{12\varepsilon^{\prime}}\right)<\frac{1}{n^{4c}}\,.$
@@
Hence, by the union bound over $n$ nodes, with probability at least $1-\frac{1}{n^{3c}}$, each good receiver satisfies its matching condition. Each hopeful, good edge has $1/4$ probability that it is going from a proposer to a receiver. The expected number of hopeful edges that are good and are going from proposers to receivers is then at least $\frac{H}{8}\cdot(1-1/n^{3})$; these are also the expected number of hopeful edges that become unhopeful. By the Markov bound, the probability that at most $1615H/$ 16 hopeful edges remain hopeful is $\frac{7/8+\frac{1}{n^{3}}}{15/16}=14/15+\frac{15}{16n^{3}}$. This means that with probability at least $1/16$ (assuming $n>10$—if $n\leq 10$, we can run our algorithm from Section 4 in $O(1)$ rounds), at least $H/16$ of the hopeful edges become unhopeful. ∎
@@
Using all of the above lemmas, we prove Lemma 5.2, which states the final utility guarantee of our algorithm: It is guaranteed to return a maximal matching in $O(\log n)$ rounds, with high probability.
Proof of Lemma 5.2.
@@
By Lemma 5.5, in each round $i$, at least $1/16$ of the hopeful edges become unhopeful with probability at least $1/16$. In $32\cdot 16c\log(n)$ rounds, the expected number of rounds for which at least $1/16$ of the hopeful edges become unhopeful is $32c\log(n)$ rounds. Because the Chernoff bound holds for binomials, by Theorem A.2, if we set $\psi=0.5$, then the probability that less than $16c\log(n)$ rounds are successful is at most $\exp\left(\frac{0.5^{2}\cdot 32c\log(n)}{3}\right)<\frac{1}{n^{8c}}$. Hence, in $O(\log n)$ rounds, all edges become unhopeful.